# BlindGuard: Safeguarding LLM-based Multi-Agent Systems under Unknown Attacks

### v1

@@ Introduction
Instead of using a binary classifier, in BlindGuard, we employ a supervised contrastive learning strategy to utilize the synthetic anomalies for model training. Our core idea is to maximize the similarity among normal agents, and minimize the similarity between normal and malicious ones. This explicit optimization creates clearer decision boundaries between the normal agents and the corrupted ones, while avoiding overfitting to specific synthetic corruption patterns. Mathematically, the supervised contrastive learning loss is defined as:
@@
The anomaly score of agent $v_{i}$ increases proportionally with its deviation from the global representation pattern of the MAS in the embedding space.
Pruning-based Remediation
@@
This combination of detection and remediation mechanisms positions BlindGuard as an unsupervised defense method for real-world MAS deployments, particularly in adversarial environments where traditional defense methods fail to adapt to evolving and unknown attack strategies. An algorithmic description of BlindGuard is in Appendix B.
Topology Method PI (CSQA) PI (MMLU) PI (GSM8k) TA (InjecAgent) MA (PosionRAG) MA (CSQA)
@@
Chain No Defense - 38.33 - 34.67 - 9.83 - 48.00 - 22.33 - 26.00
G-Safeguard 100.00 19.67 98.22 17.00 98.22 4.40 100.00 10.24 100.00 6.00 94.67 5.67
DOMINANT 42.22 28.00 53.78 24.67 67.56 8.47 88.00 14.98 64.44 14.00 27.56 35.33
PREM 51.56 26.33 48.00 26.33 62.22 8.79 89.33 15.17 61.33 16.33 57.78 17.33
TAM 26.67 57.78 49.33 25.00 51.56 8.84 61.33 30.04 50.67 17.67 53.78 18.67
BlindGuard 79.11 25.00 84.89 21.33 69.33 8.47 86.22 16.38 81.33 15.67 74.67 12.33
Tree No Defense - 33.33 - 33.33 - 10.20 - 45.05 - 20.33 - 20.33
G-Safeguard 100.00 18.33 99.11 18.33 99.11 7.80 100.00 4.76 99.56 8.00 90.67 9.00
DOMINANT 47.11 25.67 56.44 20.67 68.44 6.78 88.44 15.33 64.44 11.00 29.78 22.67
PREM 52.44 25.67 41.78 21.56 53.33 8.47 85.78 16.21 56.89 11.67 58.22 18.67
TAM 56.89 25.00 54.67 21.33 54.67 8.14 61.33 32.01 58.22 12.33 55.56 17.33
BlindGuard 75.56 20.67 79.56 20.33 59.55 8.47 85.78 12.50 76.44 10.00 77.33 13.33
Star No Defense - 46.33 - 42.33 - 12.89 - 43.57 - 23.33 - 24.67
G-Safeguard 100.00 18.33 99.11 18.00 98.22 6.10 100.00 6.87 100.00 7.33 95.11 5.67
DOMINANT 47.56 30.33 55.11 24.67 69.78 7.80 89.33 14.33 62.67 14.67 27.56 35.33
PREM 51.56 32.67 45.33 27.33 59.56 10.51 93.78 14.68 62.22 12.67 56.44 25.33
TAM 60.00 28.00 64.89 25.33 68.44 8.14 71.11 26.57 62.67 20.00 59.56 18.00
BlindGuard 82.67 24.00 85.33 21.33 70.22 6.78 93.78 12.59 85.33 12.00 74.67 12.33
Random No Defense - 35.67 - 48.33 - 14.48 - 39.78 - 26.21 - 31.33
G-Safeguard 98.22 18.67 99.56 19.33 99.11 3.79 98.22 5.14 97.70 10.34 91.56 7.67
Dominant 44.89 33.67 60.00 31.00 69.78 10.51 84.63 14.93 61.33 18.33 25.78 35.33
PREM 53.33 31.67 40.85 33.09 69.78 10.51 86.22 14.49 61.33 12.33 57.78 25.33
TAM 47.11 36.67 47.56 38.00 46.22 14.91 52.00 35.78 48.44 25.33 51.11 26.33
BlindGuard 76.89 23.67 84.00 24.67 75.56 6.44 79.56 17.69 82.22 9.67 74.67 16.00
Table 1: AUC and ASR@3 of different defense methods with GPT-4o-mini serving as the backbone LLM. Following G-safeguard (Wang et al. 2025), we consider three types of attack: Prompt injection (PI), tool attack (TA), and memory attack (MA). We showcase results after round 3 communications (ASR@3), and the additional results are placed in Appendix D.
Experiments
@@
Following G-Safeguard (Wang et al. 2025), we evaluate the defense capabilities of BlindGuard against three attack strategies: (1) direct prompt attacks using adversarial samples from CSQA (Talmor et al. 2018), MMLU (Hendrycks et al. 2021) and GSM8K (Cobbe et al. 2021); (2) tool attacks constructed from the InjecAgent dataset (Zhan et al. 2024b); and (3) memory attacks configured according to PoisonRAG (Nazary, Deldjoo, and Noia 2025) and CSQA (Talmor et al. 2018).
@@
We compare our approach with the following anomalous agent detection methods. G-Safeguard (Wang et al. 2025) is a graph-based defense framework that formulates malicious agent detection as a supervised classification task using GNNs. Note that G-Safeguard serves as an upper bound in our experiments, since it uses extra ground-truth attacked agent data for supervised model training. For unsupervised methods, we take representative GAD methods for comparisons, including: DOMINANT (Ding et al. 2019), a generation-based method; PREM (Pan et al. 2023), a contrastive learning-based method; and TAM (Qiao and Pang 2023), an affinity-driven method.
Figure 3: The overall performance of MAS on the CSQA dataset after each turn of dialogue.
@@
We evaluate the defensive capabilities of BlindGuard through comprehensive experiments spanning multiple attack types, topological structures, and LLM backbones. Following G-Safeguard, our testing framework employs three primary attack methods: direct prompt, tool attack, and memory poisoning. For network topologies, we examine four distinct MAS structures - chain, tree, star, and random - to validate generalization across communication patterns. The experiments incorporate both open-source LLMs (Qwen3-30B-A3B (Yang et al. 2025), Deepseek-v3 (Liu et al. 2024)) and commercial LLMs (GPT-4o-mini) as agent backbones. Critical performance metrics include Attack Success Rate after three communication rounds (ASR@3) and Area Under Curve (AUC) of malicious agent detection. To ensure fairness and practicality, we set a budget to identify the top three agents with the highest risk in the MAS as the predicted malicious agents. More experimental setups can be found in Appendix C.
Figure 4: ASR@3 with DeepSeek-V3 and Qwen3-30B-A3B as backbone LLMs on the CSQA and PoisonRAG datasets.
Experimental Results
@@
To investigate the generalizability of BlindGuard, we conducted additional experiments using DeepSeek-V3 and Qwen3-30B-A3B as backbone LLMs on the CSQA and PoisonRAG datasets, as shown in Figure 4 (more can be found in Appendix D). Through experiments, we make the following observations. ❹ BlindGuard obtains robust defense performance when deployed with diverse LLM and topologies. As shown in Figure 4, BlindGuard maintains robust defense performance in ASR@3 and AUC across different LLM backbones and topological structures. This stable performance confirms that BlindGuard effectively captures universal adversarial patterns rather than overfitting to specific LLM or topologies. ❺ BlindGuard successfully generalizes to different attack types on the same dataset using one universal trained model. As shown in Table 1, while supervised defense methods like G-Safeguard require training specialized models for different attack types on the same dataset (e.g., CSQA), BlindGuard achieves defense using a single model for both attack types (i.e., PI and MA). This property shows the potential of BlindGuard to be a universal defense model against different unseen attacks.
@@
To investigate the scalability of BlindGuard to larger-scale MAS, we report defense performance of PoisonRAG across systems with 20 and 50 agents, as shown in Table 2. We observe that ❻ BlindGuard consistently mitigates adversarial impact across all rounds (R1–R3) in larger-scale MAS. The scalability of BlindGuard is caused by its topology-agnostic design, where hierarchical agent encoder and corruption-guided attack detector eliminate dependencies on fixed agent numbers or interaction patterns, thereby ensuring consistent performance across diverse scales. This defense under scaling demonstrates the practicality of BlindGuard for real-world large-scale MAS.
@@
Table 2: ASR@3 on different agent numbers and rounds.
@@
Figure 5: Ablation study on PoisonRAG. NL and GL denote neighbor-level and global-level features, respectively.
To study the hierarchical agent encoder’s role in BlindGuard is quantified, we conduct an ablation study on the PoisonRAG dataset. As shown in Figure 5, we observe that ❼ anomaly detection in MAS requires a combination of both local neighborhood interactions and global system context. Removing neighborhood and global context features leads to significant performance degradation, and their combined absence causes a severe drop, highlighting the critical role of structural context beyond agent-level features. This observation shows the significance of combining information at multiple levels.
Conclusion
@@
A. Related Work
LLM-based Multi-agent System
@@ A. Related Work
Recent advances in LLM-based MAS have demonstrated remarkable capabilities in general task-solving. The performance of MAS is predominantly determined by collaboration and communication among agents with diverse roles and expertise (Tran et al. 2025), where well-structured communication can improve overall effectiveness. Modern LLM-based MAS implementations employ varied collaboration strategies to optimize performance. Sequential reasoning and debate-based role specialization (Li et al. 2024c) have proven particularly effective for knowledge-intensive tasks, while centralized planning architectures demonstrate superior performance in goal-oriented scenarios. Notable frameworks include conversational agent networks in AutoGen (Wu et al. 2024), the developer-centric platform in AgentScope (Gao et al. 2024), and phase-structured software development in ChatDev (Qian et al. 2023). The effectiveness of LLM-based MAS has been closely tied to the quality of its communication topologies. Recent research has explored MAS based on graph algorithms (Zhang et al. 2025a, b). Despite their effectiveness, these graph-based MAS topologies remain vulnerable to adversarial manipulation, where malicious agents can exploit the communication structure to inject misinformation, disrupt coordination, or compromise collective decisions.
Security of LLM-based MAS
@@
Despite the effectiveness of LLM-based MAS, this advancement has introduced novel security risks, particularly threats that exploit agent memory (Chen et al. 2024) and tool-handling mechanisms (Zhan et al. 2024a). The most severe threats target message-passing mechanisms (Zhou et al. 2025), enabling malicious attackers to implant prejudiced content. NetSafe (Yu et al. 2024) pioneers the study of network structure vulnerabilities, identifying bias propagation patterns in a multi-agent utterance graph. G-Safeguard (Wang et al. 2025) advances supervised detection of compromised agents through graph neural networks and topological remediation. A-Trust (He et al. 2025) develops attention-based trust metrics by analyzing violation patterns across six fundamental trust dimensions. While these methods can mitigate certain security threats, they heavily rely on labeled malicious agents or prior knowledge of attack patterns, which may not be available in real-world MAS deployments.
B. Algorithm
@@
C. Detailed Experimental Setup
We employ the Adam optimizer (Kingma and Ba 2014) with an initial learning rate of 0.001 and L2 regularization ($\text{weight decay}\in\{5\times 10^{-5},10^{-4},2\times 10^{-4}\}$). The learning rate is dynamically adjusted using a cosine annealing scheduler ($T_{\text{max}}=10$ cycles and $\eta_{\text{min}}=10^{-5}$) to facilitate better convergence. All models are implemented with a hidden dimension of 512 and trained on 4 NVIDIA L40 GPUs.
Figure 6: The AUC with DeepSeek-V3 and Qwen3-30B-A3B as backbone LLMs on the CSQA and PoisonRAG datasets.

### v2

@@ 3.3 Corruption-Guided Attack Detector
Instead of using a binary classifier, in BlindGuard, we employ a supervised contrastive learning strategy to utilize the synthetic anomalies for model training. Our core idea is to maximize the similarity among normal agents, and minimize the similarity between normal and malicious ones. This explicit optimization creates clearer decision boundaries between the normal agents and the corrupted ones, while avoiding overfitting to specific synthetic corruption patterns. The supervised contrastive learning paradigm achieves superior separation in the embedding space through two key mechanisms: intra-class compactness and inter-class separation Khosla et al. (2020); Zhang et al. (2025a). Mathematically, the supervised contrastive learning loss is defined as:
@@
The anomaly score of agent $v_{i}$ increases proportionally with its deviation from the global representation pattern of MAS in the embedding space.
3.4 Pruning-based Remediation
@@
This combination of detection and remediation mechanisms positions BlindGuard as an unsupervised defense method for real-world MAS deployments, particularly in adversarial environments where traditional defense methods fail to adapt to evolving and unknown attacks. An algorithmic description of BlindGuard is in Appendix B.
Topology Method PI (CSQA) PI (MMLU) PI (GSM8k) TA (InjecAgent) MA (PoisonRAG) MA (CSQA)
@@
Chain No Defense - 42.33 - 38.33 - 9.83 - 48.00 - 22.33 - 26.67
G-Safeguard 100.00 19.33 99.11 19.33 98.22 4.40 100.00 10.24 100.00 4.00 94.67 7.67
DOMINANT 47.11 30.33 59.56 24.67 67.56 8.47 88.00 14.98 64.00 11.00 36.44 18.67
PREM 50.67 29.33 44.89 25.00 62.22 8.79 89.33 15.17 61.33 6.00 68.89 15.67
TAM 55.11 27.33 52.89 23.67 51.56 8.84 61.33 30.04 52.00 14.67 51.11 13.33
BlindGuard 80.00 23.67 85.78 19.33 69.33 8.47 86.22 16.38 81.78 10.00 73.78 7.00
Tree No Defense - 33.00 - 32.00 - 10.20 - 45.05 - 18.00 - 21.33
G-Safeguard 99.56 18.67 98.67 18.33 99.11 7.80 100.00 4.76 100.00 3.00 93.78 6.67
DOMINANT 44.89 27.00 61.33 19.33 68.44 6.78 88.44 15.33 65.33 14.33 40.44 21.67
PREM 50.67 22.67 44.44 24.00 53.33 8.47 85.78 16.21 62.22 8.33 64.44 14.00
TAM 53.78 26.00 55.56 22.00 54.67 8.14 61.33 32.01 56.44 12.00 54.67 13.00
BlindGuard 74.67 24.00 83.56 20.67 59.55 8.47 85.78 12.50 76.89 7.00 72.89 7.67
Star No Defense - 50.33 - 39.00 - 12.89 - 43.57 - 26.00 - 27.00
G-Safeguard 100.00 19.33 98.67 17.33 98.22 6.10 100.00 6.87 100.00 1.67 96.44 1.67
DOMINANT 38.22 38.67 62.67 26.67 69.78 7.80 89.33 14.33 65.33 13.67 36.44 26.67
PREM 56.89 34.33 43.11 31.00 59.56 10.51 93.78 14.68 62.22 9.00 70.22 20.00
TAM 57.33 28.67 64.44 28.33 68.44 8.14 71.11 26.57 66.22 19.67 61.33 19.00
BlindGuard 85.78 22.00 86.22 20.33 70.22 6.78 93.78 12.59 89.33 10.00 76.44 9.33
Random No Defense - 42.00 - 41.00 - 14.48 - 39.78 - 24.67 - 24.00
G-Safeguard 99.11 19.00 99.11 19.00 99.11 3.79 98.22 5.14 98.67 7.00 95.56 5.00
Dominant 46.67 37.00 57.33 29.00 69.78 10.51 84.63 14.93 65.33 17.00 38.22 34.33
PREM 52.44 34.67 48.89 37.00 69.78 10.51 86.22 14.49 62.22 13.67 64.44 16.67
TAM 45.33 35.33 49.33 38.33 46.22 14.91 52.00 35.78 49.78 23.67 52.00 26.67
BlindGuard 75.11 30.00 85.33 26.67 75.56 6.44 79.56 17.69 82.67 6.67 75.56 8.67
Table 1: AUC $\left(\uparrow\right)$ and ASR@3 $\left(\downarrow\right)$ of different defense methods with GPT-4o-mini serving as the backbone LLM. Following G-Safeguard, we consider three types of attack: Prompt injection (PI), tool attack (TA), and memory attack (MA). We showcase results after round 3 communications (ASR@3), and the additional results are placed in Appendix E.
4 Experiments
@@ 4.1 Experimental Setups
Following G-Safeguard Wang et al. (2025), we evaluate the defense capabilities of BlindGuard against three attack strategies: (1) direct prompt attacks using adversarial samples from CSQA Talmor et al. (2018), MMLU Hendrycks et al. (2021) and GSM8K Cobbe et al. (2021); (2) tool attacks constructed from the InjecAgent dataset Zhan et al. (2024a); and (3) memory attacks configured according to PoisonRAG Nazary et al. (2025) and CSQA Talmor et al. (2018).
@@
We compare our method with the following anomalous agent detection methods. G-Safeguard (Wang et al., 2025) is a graph-based defense framework that formulates malicious agent detection as a supervised classification task. Note that G-Safeguard serves as an upper bound in our experiments, since it uses extra ground-truth attacked agent data for supervised model training. For unsupervised methods, we take representative GAD methods for comparisons, including: DOMINANT (Ding et al., 2019), a generation-based method; PREM Pan et al. (2023), a contrastive learning-based method; and TAM (Qiao and Pang, 2023), an affinity-driven method.
@@
We evaluate the defensive capabilities of BlindGuard through comprehensive experiments spanning multiple attack types, topological structures, and LLM backbones. Following G-Safeguard, we simulate scenarios where exactly three agents are compromised in the MAS. Our testing framework employs three primary attack methods: direct prompt, tool attack, and memory poisoning. For network topologies, we examine four distinct MAS structures - chain, tree, star, and random - to validate generalization across communication patterns. The experiments incorporate both open-source LLMs (Qwen3-30B-A3B (Yang et al., 2025), Deepseek-v3 (Liu et al., 2024)) and commercial LLMs (GPT-4o-mini) as agent backbones. Critical performance metrics include Attack Success Rate after three communication rounds (ASR@3) and Area Under Curve (AUC) of malicious agent detection. To ensure fairness and practicality, we set a budget to identify the top three agents with the highest risk in the MAS as the predicted malicious agents. More experimental setups can be found in Appendix D.
4.2 Experimental Results
@@
Figure 3: The overall performance of MAS on the CSQA (PI) dataset after each turn of dialogue.
Figure 4: ASR@3 with DeepSeek-V3 and Qwen3-30B-A3B as backbone LLMs on CSQA (PI) and PoisonRAG.
@@
To investigate the generalizability of BlindGuard, we conducted additional experiments using DeepSeek-V3 and Qwen3-30B-A3B as backbone LLMs on the CSQA and PoisonRAG datasets, as shown in Figure 4 (more can be found in Appendix E). Through experiments, we make the following observations. ❹ BlindGuard obtains robust defense performance when deployed with diverse LLM and topologies. As shown in Figure 4, BlindGuard maintains robust defense performance in ASR@3 and AUC across different LLM backbones and topological structures. This stable performance confirms that BlindGuard effectively captures universal adversarial patterns rather than overfitting to specific LLM or topologies. ❺ BlindGuard demonstrates superior generalization against unseen attacks compared to the supervised paradigm. As shown in Table 2, when the defense model trained on PI is tested against the unseen MA on the CSQA dataset, BlindGuard consistently outperforms the G-Safeguard across most topologies. This phenomenon indicates that the supervised method, reliant on known attack patterns for training, suffers from a notable performance drop when facing an unknown attack type. In contrast, BlindGuard, which learns the inherent patterns of normal agent behaviors, establishes a more generalized decision boundary. This advantage allows it to effectively identify deviations caused by diverse adversarial strategies without prior exposure, showcasing its strong generalization capability as a practical and attack-agnostic defense solution.
Topology G-Safeguard BlinGuard
ASR@3 AUC ASR@3 AUC
Chain 19.00 59.56 9.00 73.33
Tree 16.70 60.44 8.00 78.67
Star 10.67 64.00 7.33 75.56
Random 13.33 60.44 13.33 78.22
Table 2: AUC $\left(\uparrow\right)$ and ASR@3 $\left(\downarrow\right)$ on the CSQA (MA).
@@
To investigate the scalability of BlindGuard to larger-scale MAS, we report defense performance of PoisonRAG across systems with 20 and 50 agents, as shown in Table 3. We observe that ❻ BlindGuard consistently mitigates adversarial impact across all rounds (R1–R3) in larger-scale MAS. The scalability of BlindGuard is caused by its topology-agnostic design, where hierarchical agent encoder and corruption-guided attack detector eliminate dependencies on fixed agent numbers or interaction patterns, thereby ensuring consistent performance across diverse scales. This defense under scaling demonstrates the practicality of BlindGuard for large-scale MAS.
@@
Table 3: ASR@3 $\left(\downarrow\right)$ on different agent numbers.
@@
Figure 5: Ablation study on PoisonRAG. NL and GL denote neighbor-level and global-level features.
To study the hierarchical agent encoder’s role in BlindGuard is quantified, we conduct an ablation study on the PoisonRAG dataset. As shown in Figure 5, we observe that ❼ anomaly detection in MAS requires a combination of both local neighborhood interactions and global system context. Removing neighborhood and global context features leads to significant performance degradation, and their combined absence causes a severe drop, highlighting the critical role of structural context beyond agent-level features. A node may look locally consistent under colluding attackers, but still be globally misaligned with the system’s overall intent. Hence the global feature increases expressiveness by providing an additional reference axis, not by providing node-unique information directly. This observation shows the significance of combining information at multiple levels.
5 Conclusion
@@
Limitations
While BlindGuard demonstrates effective capabilities in identifying malicious agents through unsupervised graph anomaly detection, it is important to note several limitations. First, the current evaluation is limited to simulated multi-agent environments, and future work should validate the framework in more diverse and open-world scenarios to better assess its generalizability. Second, as an unsupervised detection-based approach, BlindGuard cannot preemptively prevent malicious agents from infiltrating the system, but rather mitigates adversarial propagation after intrusion has occurred. Our intent is to position BlindGuard as a complementary layer within a defense strategy, not a replacement for preventive methods. If upstream methods are evaded, BlindGuard runs at the boundary, immediately after the first generation and before the next propagation. Therefore, developing more adaptive defense strategies remains an important direction for future research.
Ethical Considerations
Our research involves no human subjects, animal experiments, or sensitive data. All work is based on synthetic or publicly available data in simulated environments. We foresee no ethical risks or conflicts of interest. We are committed to maintaining the highest standards of research integrity to ensure full compliance with ethical guidelines.
Acknowledgments
This work was supported by a grant from the National Natural Science Foundation of China under grants (No.62372211), the Science and Technology Development Program of Jilin Province (No.20250102216JC). The work of Y. Liu was partially supported by the Australian Research Council (ARC) under Grant No.DE260101172.
Appendix A Related Work
A.1 LLM-based Multi-agent System
@@
Recent advances in LLM-based MAS have demonstrated remarkable capabilities in general task-solving. The performance of MAS is predominantly determined by collaboration and communication among agents with diverse roles and expertise (Tran et al., 2025). Modern LLM-based MAS implementations employ varied collaboration strategies to optimize performance (Li et al., 2026b). Sequential reasoning and debate-based role specialization (Li et al., 2024c) have proven particularly effective for knowledge-intensive tasks, while centralized planning architectures demonstrate superior performance in goal-oriented scenarios. Notable frameworks include conversational agent networks in AutoGen (Wu et al., 2024), the developer-centric platform in AgentScope (Gao et al., 2024), and phase-structured software development in ChatDev (Qian et al., 2023). Recent research has explored MAS based on graph algorithms (Zhang et al., 2025b, c; Shen et al., 2025a; Li et al., 2026a). Despite their effectiveness, these graph-based MAS topologies remain vulnerable to adversarial manipulation, where malicious agents can exploit the communication structure to inject misinformation, disrupt coordination, or compromise collective decisions.
A.2 Security of LLM-based MAS
@@
Despite the effectiveness of LLM-based MAS, this advancement has introduced novel security risks, particularly threats that exploit agent memory (Chen et al., 2024) and tool-handling mechanisms (Zhan et al., 2024b). Furthermore, the inherent vulnerabilities of LLMs themselves, such as hallucinations and unreliable reasoning (Qian et al., 2026; Bi et al., 2025; Huang et al., 2025; Yuan et al., 2026; Shen et al., 2025b; Zhang et al., 2025d, 2026; Wang et al., 2026), can be further amplified in multi-agent settings (He et al., 2025; Shen et al., 2025c). The most severe threats target message-passing mechanisms (Zhou et al., 2025; Pan et al., 2026a), enabling malicious attackers to implant prejudiced content. NetSafe (Yu et al., 2024) pioneers the study of network structure vulnerabilities, identifying bias propagation patterns in a multi-agent utterance graph. G-Safeguard (Wang et al., 2025) advances supervised detection of compromised agents through graph neural networks and topological remediation. While these methods can mitigate certain security threats, they heavily rely on labeled malicious agents or prior knowledge of attack patterns, which may not be available in real-world MAS deployments.
Appendix B Algorithm
@@
Figure 6: The AUC with DeepSeek-V3 and Qwen3-30B-A3B as backbone LLMs on the CSQA and PoisonRAG.